OpenStudy (anonymous):

FAN+MEDAL!! Quadrilateral OPQR is inscribed inside a circle as shown below. What is the measure of angle R? You must show all work and calculations to receive credit. Circle N is shown with an inscribed quadrilateral labeled OPQR. O is labeled 2x degrees. P is labeled y degrees. Q is labeled 2x plus 4 degrees. R is labeled 3y plus 8 degrees. Please help I don't umderstand..

OpenStudy (anonymous):

anyone?

OpenStudy (anonymous):

anyone going to help me plz??

Vocaloid (vocaloid):

opposite angles in an inscribed quadrilateral sum to 180 degrees, so... 2x + (2x + 4) = 180 (3y + 8) + y = 180 solve for x and y

OpenStudy (anonymous):

oh okay thank you, anything else after that??

Vocaloid (vocaloid):

um, there's nothing more to tell you, all you need to do is solve the two equations

OpenStudy (anonymous):

k workin on it now

Vocaloid (vocaloid):

well, actually, you can skip the first equation and just solve the second one

Vocaloid (vocaloid):

(3y + 8) + y = 180 find y

OpenStudy (anonymous):

dang it

OpenStudy (anonymous):

y=43 is that right?

Vocaloid (vocaloid):

yes, y = 43 now, the question wants the value of angle R, and angle R = 3y + 8, so R = ?

OpenStudy (anonymous):

k workin on it...

OpenStudy (anonymous):

how do u do it?? im having a total brain fart right now... hmm

OpenStudy (anonymous):

omg i feel so stupid

Vocaloid (vocaloid):

y = 43 3y + 8 = 3(43) + 8 = ?

OpenStudy (anonymous):

oh!! okay thx

OpenStudy (anonymous):

it gives me 43 again..

Vocaloid (vocaloid):

3(43) = ?

OpenStudy (anonymous):

129

Vocaloid (vocaloid):

so 3(43) + 8 = ?

OpenStudy (anonymous):

137

Vocaloid (vocaloid):

yup, there's our final answer

OpenStudy (anonymous):

thats it?? wow that was so much eaiser thn i thought

OpenStudy (anonymous):

thank you very much for helping me out! i really appreciate it!